a=I+2j-3k

b=4i+j-7k.

Find the modulus and direction cosine of the following

b+a

b-a







1
Expert's answer
2021-06-23T17:05:26-0400

1)


"\bold{b+a}=(5,3,-10)\\b+a=\sqrt{134}\\
(\cos a,\cos b,\cos c)=\frac{1}{\sqrt{134}}(5,3,-10)"

2)


"\bold{b-a}=(3,-1,-4)\\b-a=\sqrt{26}\\
(\cos a,\cos b,\cos c)=\frac{1}{\sqrt{26}}(3,-1,-4)"